OpenStudy (anonymous):

f(x)=x2+ 1x + 2 ___ 4 g(x)=x^3+ 1 ___x^2+X 2 Find (f + g)(a).

OpenStudy (anonymous):

simply add the two together in order to do this tho first you have to get the denominators to equal

OpenStudy (anonymous):

then replace x with a

OpenStudy (anonymous):

? i'm horrible at math, i can't even do that much

OpenStudy (anonymous):

times the bottom by 1 but using 2/2

OpenStudy (anonymous):

so your bottom equation is \[\frac{x^3+x^2+x}{2}\] correct?

OpenStudy (anonymous):

if so multiply by 2/2 and you'll get \[\frac{2x^3+2x^2+2x}{4}\]

OpenStudy (anonymous):

now you can add them since their denominators are the same

OpenStudy (anonymous):

\[(f+g)=\frac{x^2+x+2}{4}+\frac{2x^3+2x^2+2x}{4}=\frac{3x^3+3x^2+2x+2}{4}\]

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

now simply interchange x with a

OpenStudy (anonymous):

\[\frac{3a^3+3a^2+2a+2}{4}=\frac{1}{4}(3a^3+3a^2+2a+2)\]

OpenStudy (anonymous):

since a has no given value that will be your answer

OpenStudy (anonymous):

ok, i keep looking it over and over to see if i understand and i am slowly getting it... thank you!.. i'm going to keep looking at it so i understand it better.